\begin{thebibliography}{99}

%1\cite{Tanabashi:2018oca}
\bibitem{Tanabashi:2018oca}
  M.~Tanabashi {\it et al.} [Particle Data Group],
  Review of Particle Physics,
  Phys.\ Rev.\ D {\bf 98},  030001 (2018).
%  doi:10.1103/PhysRevD.98.030001
  %%CITATION = doi:10.1103/PhysRevD.98.030001;%%
  %2188 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Chen:2016spr}
\bibitem{Chen:2016spr}
  H.~X.~Chen, W.~Chen, X.~Liu, Y.~R.~Liu, and S.~L.~Zhu,
  A review of the open charm and open bottom systems,
  Rept.\ Prog.\ Phys.\  {\bf 80}, 076201 (2017).
%  doi:10.1088/1361-6633/aa6420
%  [arXiv:1609.08928 [hep-ph]].
  %%CITATION = doi:10.1088/1361-6633/aa6420;%%
  %121 citations counted in INSPIRE as of 20 Aug 2019



%\cite{Cheng:2015iom}
\bibitem{Cheng:2015iom}
  H.~Y.~Cheng,
  Charmed baryons circa 2015,
  Front.\ Phys.\ (Beijing) {\bf 10},  101406 (2015).
%  doi:10.1007/s11467-015-0483-z
  %%CITATION = doi:10.1007/s11467-015-0483-z;%%
  %35 citations counted in INSPIRE as of 20 Aug 2019

%\cite{Crede:2013sze}
\bibitem{Crede:2013sze}
  V.~Crede and W.~Roberts,
  Progress towards understanding baryon resonances,
  Rept.\ Prog.\ Phys.\  {\bf 76}, 076301 (2013).
%  doi:10.1088/0034-4885/76/7/076301
%  [arXiv:1302.7299 [nucl-ex]].
  %%CITATION = doi:10.1088/0034-4885/76/7/076301;%%
  %141 citations counted in INSPIRE as of 20 Aug 2019


%\cite{Klempt:2009pi}
\bibitem{Klempt:2009pi}
  E.~Klempt and J.~M.~Richard,
  Baryon spectroscopy,
  Rev.\ Mod.\ Phys.\  {\bf 82}, 1095 (2010).
%  doi:10.1103/RevModPhys.82.1095
%  [arXiv:0901.2055 [hep-ph]].
  %%CITATION = doi:10.1103/RevModPhys.82.1095;%%
  %293 citations counted in INSPIRE as of 20 Aug 2019




%2\cite{Aaij:2017nav}
\bibitem{Aaij:2017nav}
  R.~Aaij {\it et al.} [LHCb Collaboration],
  Observation of five new narrow $\Omega_c^0$ states decaying to $\Xi_c^+ K^-$,
  Phys.\ Rev.\ Lett.\  {\bf 118},  182001 (2017).
%  doi:10.1103/PhysRevLett.118.182001
%  [arXiv:1703.04639 [hep-ex]].
  %%CITATION = doi:10.1103/PhysRevLett.118.182001;%%
  %144 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Cheng:2017ove}
\bibitem{Cheng:2017ove}
  H.~Y.~Cheng and C.~W.~Chiang,
  Quantum numbers of $\Omega_c$ states and other charmed baryons,
  Phys.\ Rev.\ D {\bf 95},  094018 (2017).
%  doi:10.1103/PhysRevD.95.094018
%%  [arXiv:1704.00396 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.094018;%%
  %34 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Chen:2017gnu}
\bibitem{Chen:2017gnu}
  B.~Chen and X.~Liu,
  New $\Omega_c^0$ baryons discovered by LHCb as the members of $1P$ and $2S$ states,
  Phys.\ Rev.\ D {\bf 96},  094015 (2017).
%  doi:10.1103/PhysRevD.96.094015
%  [arXiv:1704.02583 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.094015;%%
  %35 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Padmanath:2017lng}
\bibitem{Padmanath:2017lng}
  M.~Padmanath and N.~Mathur,
  Quantum Numbers of Recently Discovered $\Omega^{0}_{c}$ Baryons from Lattice QCD,
  Phys.\ Rev.\ Lett.\  {\bf 119},  042001 (2017).
%  doi:10.1103/PhysRevLett.119.042001
%  [arXiv:1704.00259 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevLett.119.042001;%%
  %41 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Wang:2017vnc}
\bibitem{Wang:2017vnc}
  W.~Wang and R.~L.~Zhu,
  Interpretation of the newly observed $\Omega_c^0$ resonances,
  Phys.\ Rev.\ D {\bf 96},  014024 (2017).
%  doi:10.1103/PhysRevD.96.014024
%  [arXiv:1704.00179 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.014024;%%
  %46 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Karliner:2017kfm}
\bibitem{Karliner:2017kfm}
  M.~Karliner and J.~L.~Rosner,
  Very narrow excited $\Omega_c$ baryons,
  Phys.\ Rev.\ D {\bf 95},  114012 (2017).
%  doi:10.1103/PhysRevD.95.114012
%  [arXiv:1703.07774 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.114012;%%
  %55 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Chen:2017sci}
\bibitem{Chen:2017sci}
  H.~X.~Chen, Q.~Mao, W.~Chen, A.~Hosaka, X.~Liu, and S.~L.~Zhu,
  Decay properties of $P$-wave charmed baryons from light-cone QCD sum rules,
  Phys.\ Rev.\ D {\bf 95},  094008 (2017).
%  doi:10.1103/PhysRevD.95.094008
%  [arXiv:1703.07703 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.094008;%%
  %46 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Agaev:2017lip}
\bibitem{Agaev:2017lip}
  S.~S.~Agaev, K.~Azizi, and H.~Sundu,
  Interpretation of the new $\Omega_c^{0}$ states via their mass and width,
  Eur.\ Phys.\ J.\ C {\bf 77},  395 (2017).
%  doi:10.1140/epjc/s10052-017-4953-z
%  [arXiv:1704.04928 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-017-4953-z;%%
  %27 citations counted in INSPIRE as of 06 Aug 2019

%\cite{Wang:2017zjw}
\bibitem{Wang:2017zjw}
  Z.~G.~Wang,
  Analysis of $\Omega_c(3000)$, $\Omega_c(3050)$, $\Omega_c(3066)$, $\Omega_c(3090)$ and $\Omega_c(3119)$ with QCD sum rules,
  Eur.\ Phys.\ J.\ C {\bf 77}, 325 (2017).
%  doi:10.1140/epjc/s10052-017-4895-5
%  [arXiv:1704.01854 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-017-4895-5;%%
  %31 citations counted in INSPIRE as of 20 Aug 2019



%\cite{Wang:2017hej}
\bibitem{Wang:2017hej}
  K.~L.~Wang, L.~Y.~Xiao, X.~H.~Zhong, and Q.~Zhao,
  Understanding the newly observed $\Omega_c$ states through their decays,
  Phys.\ Rev.\ D {\bf 95},  116010 (2017).
%  doi:10.1103/PhysRevD.95.116010
%  [arXiv:1703.09130 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.116010;%%
  %43 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Aaij:2018yqz}
\bibitem{Aaij:2018yqz}
  R.~Aaij {\it et al.} [LHCb Collaboration],
  Observation of a new $\Xi_b^-$ resonance,
  Phys.\ Rev.\ Lett.\  {\bf 121},  072002 (2018).
%  doi:10.1103/PhysRevLett.121.072002
%  [arXiv:1805.09418 [hep-ex]].
  %%CITATION = doi:10.1103/PhysRevLett.121.072002;%%
  %30 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Aaij:2018tnn}
\bibitem{Aaij:2018tnn}
  R.~Aaij {\it et al.} [LHCb Collaboration],
  Observation of Two Resonances in the $\Lambda_b^0 \pi^\pm$ Systems and Precise Measurement of $\Sigma_b^\pm$ and $\Sigma_b^{*\pm}$ properties,
  Phys.\ Rev.\ Lett.\  {\bf 122},  012001 (2019).
%  doi:10.1103/PhysRevLett.122.012001
%  [arXiv:1809.07752 [hep-ex]].
  %%CITATION = doi:10.1103/PhysRevLett.122.012001;%%
  %17 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Yang:2018lzg}
\bibitem{Yang:2018lzg}
  P.~Yang, J.~J.~Guo, and A.~Zhang,
  Identification of the newly observed $\Sigma_b(6097)^\pm$ baryons from their strong decays,
  Phys.\ Rev.\ D {\bf 99},  034018 (2019).
%  doi:10.1103/PhysRevD.99.034018
%  [arXiv:1810.06947 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.99.034018;%%
  %4 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Aliev:2018vye}
\bibitem{Aliev:2018vye}
  T.~M.~Aliev, K.~Azizi, Y.~Sarac, and H.~Sundu,
  Determination of the quantum numbers of $\Sigma_b(6097)^{\pm}$ via their strong decays,
  Phys.\ Rev.\ D {\bf 99},  094003 (2019).
%  doi:10.1103/PhysRevD.99.094003
%  [arXiv:1811.05686 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.99.094003;%%
  %2 citations counted in INSPIRE as of 06 Aug 2019

 %\cite{Cui:2019dzj}
\bibitem{Cui:2019dzj}
  E.~L.~Cui, H.~M.~Yang, H.~X.~Chen, and A.~Hosaka,
  Identifying the $\Xi_{b}(6227)$ and $\Sigma_{b}(6097)$ as $P$-wave bottom baryons of $J^P = 3/2^-$,
  Phys.\ Rev.\ D {\bf 99}, 094021 (2019).
%  doi:10.1103/PhysRevD.99.094021
%  [arXiv:1903.10369 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.99.094021;%%
  %1 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Wang:2018fjm}
\bibitem{Wang:2018fjm}
  K.~L.~Wang, Q.~F.~L\"u, and X.~H.~Zhong,
  Interpretation of the newly observed $\Sigma_b(6097)^{\pm}$ and $\Xi_b(6227)^-$ states as the $P$-wave bottom baryons,
  Phys.\ Rev.\ D {\bf 99},  014011 (2019).
%  doi:10.1103/PhysRevD.99.014011
%  [arXiv:1810.02205 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.99.014011;%%
  %7 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Chen:2018orb}
\bibitem{Chen:2018orb}
  B.~Chen, K.~W.~Wei, X.~Liu, and A.~Zhang,
  Role of newly discovered $\Xi_b(6227)^-$ for constructing excited bottom baryon family,
  Phys.\ Rev.\ D {\bf 98}, 031502 (2018).
%  doi:10.1103/PhysRevD.98.031502
%  [arXiv:1805.10826 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.98.031502;%%
  %11 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Chen:2018vuc}
\bibitem{Chen:2018vuc}
  B.~Chen and X.~Liu,
  Assigning the newly reported $\Sigma_b(6097)$ as a $P$-wave excited state and predicting its partners,
  Phys.\ Rev.\ D {\bf 98}, 074032 (2018).
%  doi:10.1103/PhysRevD.98.074032
%  [arXiv:1810.00389 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.98.074032;%%
  %6 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Aaij:2019amv}
\bibitem{Aaij:2019amv}
  R.~Aaij {\it et al.} [LHCb Collaboration],
  Observation of new resonances in the $\Lambda_{b}^0\pi^+\pi^-$ system,
  arXiv:1907.13598 [hep-ex].
  %%CITATION = ARXIV:1907.13598;%%
  %1 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Capstick:1986bm}
\bibitem{Capstick:1986bm}
  S.~Capstick and N.~Isgur,
  Baryons in a Relativized Quark Model with Chromodynamics,
  Phys.\ Rev.\ D {\bf 34}, 2809 (1986).
%  [AIP Conf.\ Proc.\  {\bf 132}, 267 (1985)].
%  doi:10.1103/PhysRevD.34.2809, 10.1063/1.35361
  %%CITATION = doi:10.1103/PhysRevD.34.2809, 10.1063/1.35361;%%
  %1183 citations counted in INSPIRE as of 06 Aug 2019

 %\cite{Ebert:2007nw}
\bibitem{Ebert:2007nw}
  D.~Ebert, R.~N.~Faustov, and V.~O.~Galkin,
  Masses of excited heavy baryons in the relativistic quark model,
  Phys.\ Lett.\ B {\bf 659}, 612 (2008).
%  doi:10.1016/j.physletb.2007.11.037
%  [arXiv:0705.2957 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2007.11.037;%%
  %157 citations counted in INSPIRE as of 06 Aug 2019


 %\cite{Roberts:2007ni}
\bibitem{Roberts:2007ni}
  W.~Roberts and M.~Pervin,
  Heavy baryons in a quark model,
  Int.\ J.\ Mod.\ Phys.\ A {\bf 23}, 2817 (2008).
%  doi:10.1142/S0217751X08041219
%  [arXiv:0711.2492 [nucl-th]].
  %%CITATION = doi:10.1142/S0217751X08041219;%%
  %264 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Ebert:2011kk}
\bibitem{Ebert:2011kk}
  D.~Ebert, R.~N.~Faustov, and V.~O.~Galkin,
  Spectroscopy and Regge trajectories of heavy baryons in the relativistic quark-diquark picture,
  Phys.\ Rev.\ D {\bf 84}, 014025 (2011).
%  doi:10.1103/PhysRevD.84.014025
%  [arXiv:1105.0583 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.84.014025;%%
  %124 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Mao:2015gya}
\bibitem{Mao:2015gya}
  Q.~Mao, H.~X.~Chen, W.~Chen, A.~Hosaka, X.~Liu, and S.~L.~Zhu,
  QCD sum rule calculation for P-wave bottom baryons,
  Phys.\ Rev.\ D {\bf 92}, 114007 (2015).
%  doi:10.1103/PhysRevD.92.114007
%  [arXiv:1510.05267 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.92.114007;%%
  %35 citations counted in INSPIRE as of 06 Aug 2019

 %\cite{Chen:2014nyo}
\bibitem{Chen:2014nyo}
  B.~Chen, K.~W.~Wei, and A.~Zhang,
  Assignments of $\Lambda_Q$ and $\Xi_Q$ baryons in the heavy quark-light diquark picture,
  Eur.\ Phys.\ J.\ A {\bf 51}, 82 (2015).
%  doi:10.1140/epja/i2015-15082-3
%  [arXiv:1406.6561 [hep-ph]].
  %%CITATION = doi:10.1140/epja/i2015-15082-3;%%
  %48 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Garcilazo:2007eh}
\bibitem{Garcilazo:2007eh}
  H.~Garcilazo, J.~Vijande, and A.~Valcarce,
  Faddeev study of heavy baryon spectroscopy,
  J.\ Phys.\ G {\bf 34}, 961 (2007).
%  doi:10.1088/0954-3899/34/5/014
%  [hep-ph/0703257].
  %%CITATION = doi:10.1088/0954-3899/34/5/014;%%
  %101 citations counted in INSPIRE as of 06 Aug 2019


 %\cite{Karliner:2015ema}
\bibitem{Karliner:2015ema}
  M.~Karliner and J.~L.~Rosner,
  Prospects for observing the lowest-lying odd-parity $\Sigma_c$ and $\Sigma_b$ baryons,
  Phys.\ Rev.\ D {\bf 92}, 074026 (2015).
%  doi:10.1103/PhysRevD.92.074026
%  [arXiv:1506.01702 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.92.074026;%%
  %12 citations counted in INSPIRE as of 06 Aug 2019









 %\cite{Thakkar:2016dna}
\bibitem{Thakkar:2016dna}
  K.~Thakkar, Z.~Shah, A.~K.~Rai, and P.~C.~Vinodkumar,
  Excited State Mass spectra and Regge trajectories of Bottom Baryons,
  Nucl.\ Phys.\ A {\bf 965}, 57 (2017).
%  doi:10.1016/j.nuclphysa.2017.05.087
%  [arXiv:1610.00411 [nucl-th]].
  %%CITATION = doi:10.1016/j.nuclphysa.2017.05.087;%%
  %26 citations counted in INSPIRE as of 06 Aug 2019





%\cite{Wang:2017kfr}
\bibitem{Wang:2017kfr}
  K.~L.~Wang, Y.~X.~Yao, X.~H.~Zhong, and Q.~Zhao,
  Strong and radiative decays of the low-lying $S$- and $P$-wave singly heavy baryons,
  Phys.\ Rev.\ D {\bf 96},  116016 (2017).
%  doi:10.1103/PhysRevD.96.116016
%  [arXiv:1709.04268 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.116016;%%
  %27 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Capstick:2000qj}
\bibitem{Capstick:2000qj}
  S.~Capstick and W.~Roberts,
  Quark models of baryon masses and decays,
  Prog.\ Part.\ Nucl.\ Phys.\  {\bf 45}, S241 (2000).
%  doi:10.1016/S0146-6410(00)00109-5
%  [nucl-th/0008028].
  %%CITATION = doi:10.1016/S0146-6410(00)00109-5;%%
  %345 citations counted in INSPIRE as of 06 Aug 2019


 %\cite{Ebert:2005xj}
\bibitem{Ebert:2005xj}
  D.~Ebert, R.~N.~Faustov, and V.~O.~Galkin,
  Masses of heavy baryons in the relativistic quark model,
  Phys.\ Rev.\ D {\bf 72}, 034026 (2005).
%  doi:10.1103/PhysRevD.72.034026
%  [hep-ph/0504112].
  %%CITATION = doi:10.1103/PhysRevD.72.034026;%%
  %163 citations counted in INSPIRE as of 06 Aug 2019

%\cite{Yoshida:2015tia}
\bibitem{Yoshida:2015tia}
  T.~Yoshida, E.~Hiyama, A.~Hosaka, M.~Oka, and K.~Sadato,
  Spectrum of heavy baryons in the quark model,
  Phys.\ Rev.\ D {\bf 92},  114029 (2015).
%  doi:10.1103/PhysRevD.92.114029
%  [arXiv:1510.01067 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.92.114029;%%
  %73 citations counted in INSPIRE as of 06 Aug 2019


%\cite{Valcarce:2008dr}
\bibitem{Valcarce:2008dr}
  A.~Valcarce, H.~Garcilazo, and J.~Vijande,
  Towards an understanding of heavy baryon spectroscopy,
  Eur.\ Phys.\ J.\ A {\bf 37}, 217 (2008).
%  doi:10.1140/epja/i2008-10616-4
%  [arXiv:0807.2973 [hep-ph]].
  %%CITATION = doi:10.1140/epja/i2008-10616-4;%%
  %103 citations counted in INSPIRE as of 06 Aug 2019



%\cite{Karliner:2008sv}
\bibitem{Karliner:2008sv}
  M.~Karliner, B.~Keren-Zur, H.~J.~Lipkin, and J.~L.~Rosner,
  The Quark Model and $b$ Baryons,
  Annals Phys.\  {\bf 324}, 2 (2009).
%  doi:10.1016/j.aop.2008.05.003
%  [arXiv:0804.1575 [hep-ph]].
  %%CITATION = doi:10.1016/j.aop.2008.05.003;%%
  %109 citations counted in INSPIRE as of 06 Aug 2019

%\cite{Chen:2016phw}
\bibitem{Chen:2016phw}
  H.~X.~Chen, Q.~Mao, A.~Hosaka, X.~Liu and S.~L.~Zhu,
  D-wave charmed and bottomed baryons from QCD sum rules,
  Phys.\ Rev.\ D {\bf 94},  114016 (2016).
%  doi:10.1103/PhysRevD.94.114016
%  [arXiv:1611.02677 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.94.114016;%%
  %37 citations counted in INSPIRE as of 20 Aug 2019



%\cite{Liang:2019aag}
\bibitem{Liang:2019aag}
  W.~Liang, Q.~F.~L\"u, and X.~H.~Zhong,
  Canonical interpretation of the newly observed $\Lambda_b(6146)^0$ and $\Lambda_b(6152)^0$ via strong decay behaviors,
  arXiv:1908.00223 [hep-ph].
  %%CITATION = ARXIV:1908.00223;%%






%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%  chiral quark model
%26\cite{Manohar:1983md}


%\cite{Yao:2018jmc}
\bibitem{Yao:2018jmc}
  Y.~X.~Yao, K.~L.~Wang, and X.~H.~Zhong,
  Strong and radiative decays of the low-lying $D$-wave singly heavy baryons,
  Phys.\ Rev.\ D {\bf 98},  076015 (2018).
%  doi:10.1103/PhysRevD.98.076015
%  [arXiv:1803.00364 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.98.076015;%%
  %5 citations counted in INSPIRE as of 10 Aug 2019



\bibitem{Manohar:1983md}
  A.~Manohar and H.~Georgi,
  Chiral quarks and the nonrelativistic quark Model,
  Nucl.\ Phys.\ B {\bf 234}, 189 (1984).
 % doi:10.1016/0550-3213(84)90231-1
  %%CITATION = doi:10.1016/0550-3213(84)90231-1;%%
  %1859 citations counted in INSPIRE as of 27 Mar 2017

%27\cite{Zhong:2009sk}
\bibitem{Zhong:2009sk}
  X.~H.~Zhong and Q.~Zhao,
  Strong decays of newly observed $D_{sJ}$ states in a constituent quark
  model with effective Lagrangians,
  Phys.\ Rev.\  D {\bf 81}, 014031 (2010).
  %[arXiv:0911.1856 [hep-ph]].
  %%CITATION = PHRVA,D81,014031;%%

%28\cite{Zhong:2008kd}
\bibitem{Zhong:2008kd}
  X.~H.~Zhong and Q.~Zhao,
  Strong decays of heavy-light mesons in a chiral quark model,
  Phys.\ Rev.\  D {\bf 78}, 014029 (2008).
  %[arXiv:0803.2102 [hep-ph]].
  %%CITATION = PHRVA,D78,014029;%%

%29\cite{Zhong:2010vq}
\bibitem{Zhong:2010vq}
  X.~H.~Zhong,
  Strong decays of the newly observed $D(2550)$, $D(2600)$, $D(2750)$, and $D(2760)$,
  Phys.\ Rev.\ D {\bf 82}, 114014 (2010).
 % doi:10.1103/PhysRevD.82.114014
 % [arXiv:1009.0359 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.82.114014;%%
  %36 citations counted in INSPIRE as of 26 Mar 2017

%30\cite{Xiao:2014ura}
\bibitem{Xiao:2014ura}
  L.~Y.~Xiao and X.~H.~Zhong,
  Strong decays of higher excited heavy-light mesons in a chiral quark model,
  Phys.\ Rev.\ D {\bf 90}, 074029 (2014).
%  doi:10.1103/PhysRevD.90.074029
%  [arXiv:1407.7408 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.90.074029;%%
  %12 citations counted in INSPIRE as of 26 Mar 2017

%31\cite{Zhong:2007gp}
\bibitem{Zhong:2007gp}
  X.~H.~Zhong and Q.~Zhao,
  Charmed baryon strong decays in a chiral quark model,
  Phys.\ Rev.\  D {\bf 77}, 074008 (2008).
  %%CITATION = ARXIV:0711.4645;%%

%32\cite{Liu:2012sj}
\bibitem{Liu:2012sj}
  L.~H.~Liu, L.~Y.~Xiao, and X.~H.~Zhong,
  Charm-strange baryon strong decays in a chiral quark model,
  Phys.\ Rev.\ D {\bf 86}, 034024 (2012).
 % doi:10.1103/PhysRevD.86.034024
 % [arXiv:1205.2943 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.86.034024;%%
  %18 citations counted in INSPIRE as of 26 Mar 2017



%34\cite{Xiao:2013xi}
\bibitem{Xiao:2013xi}
  L.~Y.~Xiao and X.~H.~Zhong,
  $\Xi$ baryon strong decays in a chiral quark model,
  Phys.\ Rev.\ D {\bf 87}, 094002 (2013).
 % doi:10.1103/PhysRevD.87.094002
 % [arXiv:1302.0079 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.87.094002;%%
  %9 citations counted in INSPIRE as of 26 Mar 2017

%35\cite{Xiao:2017udy}
\bibitem{Xiao:2017udy}
  L.~Y.~Xiao, K.~L.~Wang, Q.~F.~L\"{u}, X.~H.~Zhong and S.~L.~Zhu,
  Strong and radiative decays of the doubly charmed baryons,
  Phys.\ Rev.\ D {\bf 96}, 094005 (2017).
  %%CITATION = ARXIV:1708.04384;%%
  %1 citations counted in INSPIRE as of 21 Aug 2017

%36\cite{Nagahiro:2016nsx}
\bibitem{Nagahiro:2016nsx}
  H.~Nagahiro, S.~Yasui, A.~Hosaka, M.~Oka, and H.~Noumi,
  Structure of charmed baryons studied by pionic decays,
  Phys.\ Rev.\ D {\bf 95}, 014023 (2017).
 % doi:10.1103/PhysRevD.95.014023
 % [arXiv:1609.01085 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.014023;%%
  %8 citations counted in INSPIRE as of 27 Mar 2017




%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%\cite{Chen:2016iyi}
\bibitem{Chen:2016iyi}
  B.~Chen, K.~W.~Wei, X.~Liu and T.~Matsuki,
  Low-lying charmed and charmed-strange baryon states,
  Eur.\ Phys.\ J.\ C {\bf 77},  154 (2017).
 % doi:10.1140/epjc/s10052-017-4708-x
 % [arXiv:1609.07967 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-017-4708-x;%%
  %21 citations counted in INSPIRE as of 12 Aug 2019

%\cite{Chen:2017aqm}
\bibitem{Chen:2017aqm}
  B.~Chen, X.~Liu and A.~Zhang,
  Newly observed $\Lambda_c(2860)^+$ at LHCb and its \emph{D}-wave partners $\Lambda_c(2880)^+$, $\Xi_c(3055)^+$ and $\Xi_c(3080)^+$,
  Phys.\ Rev.\ D {\bf 95},  074022 (2017).
%  doi:10.1103/PhysRevD.95.074022
%  [arXiv:1702.04106 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.074022;%%
  %8 citations counted in INSPIRE as of 12 Aug 2019
















\end{thebibliography}

\documentclass[aps,twocolumn,showpacs,preprintnumbers,floats]{revtex4}\def\@cite#1#2{\textsuperscript{[{#1\if@tempswa , #2\fi}]}}

\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{color}
\usepackage{subfigure}
\usepackage{epsfig}
\usepackage{morefloats}
\usepackage{multirow}
\usepackage{graphicx,booktabs}
\usepackage{mathrsfs}
\usepackage{txfonts}
\usepackage{indentfirst}
\usepackage{graphicx,booktabs}

\usepackage{longtable,lscape}

\newcommand{\vsig}{\mbox{\boldmath$\sigma$\unboldmath}}

\usepackage{comment}

\begin{document}

\title{Interpretation of the newly observed $\Lambda_b(6146)$ and $\Lambda_b(6152)$ states in a chiral quark model}
\author{
Kai-Lei Wang$^{1}$~\footnote {E-mail: wangkaileicz@foxmail.com}, Qi-Fang L\"{u}$^{2,3}$~\footnote {E-mail: lvqifang@hunnu.edu.cn}, and Xian-Hui Zhong$^{2,3}$~\footnote {E-mail: zhongxh@hunnu.edu.cn}}

\affiliation{ 1) Department
of Electronic Information and Physics, Changzhi University, Changzhi, Shanxi,046011,China}

\affiliation{ 2) Department
of Physics, Hunan Normal University, and Key Laboratory of
Low-Dimensional Quantum Structures and Quantum Control of Ministry
of Education, Changsha 410081, China }

\affiliation{ 3) Synergetic Innovation Center for Quantum Effects and Applications (SICQEA),
Hunan Normal University, Changsha 410081, China}

\begin{abstract}
The strong decays of the low-lying $\lambda$-mode $\Lambda_b(1D,2S)$ and $\Sigma_b(2S)$ states are studied in a chiral quark model. We find that: (i) the newly observed $\Lambda_b(6146/6152)$ resonances in the $\Lambda_b\pi^+\pi^-$ spectrum  by the LHCb Collaboration
might be explained with the $\lambda$-mode $\Lambda_b(1D)$ states in the quark model. It should be emphasized that whether the structure in the $\Lambda_b\pi^+\pi^-$ spectrum correspond to two states or one state should be further clarified with more observations in future experiments. (ii) The $\Lambda_b(2S)$ $|J^P = \frac{1}{2}^+,0\rangle$ state mainly decays into $\Sigma_b^*\pi$ channel, which may be an ideal channel for searching for this $\Lambda_b(2S)$ state in future experiments. (iii) The $\Sigma_b(2S)$ $|J^P = \frac{1}{2}^+,1 \rangle$ and $|J^P = \frac{3}{2}^+,1 \rangle$ dominantly decay into $\Lambda_b\pi$ with $\Gamma \simeq 3.82$ MeV and $\Gamma \simeq 4.72$ MeV, respectively.
\end{abstract}

\maketitle

\section{Introduction}{\label{introduction}}

Understanding the heavy baryon spectrum and searching for the missing heavy baryons and new exotic states are interesting topics in hadronic physics~\cite{Tanabashi:2018oca,Chen:2016spr,Cheng:2015iom,Crede:2013sze,Klempt:2009pi}. The LHC facility provides good opportunities for us to establish the heavy baryon spectra. In 2017, five extremely narrow $\Omega_c$ states, $\Omega_c(3000)$, $\Omega_c(3050)$, $\Omega_c(3066)$, $\Omega_c(3090)$, and $\Omega_c(3119)$,
were observed  in the $\Xi_c^{+}K^-$ channel by the LHCb Collaboration~\cite{Aaij:2017nav}.
Most of them may be interpreted as the $P$-wave excited states of $\Omega_c$~\cite{Wang:2017hej,Cheng:2017ove,
Chen:2017gnu,Padmanath:2017lng,Wang:2017vnc, Karliner:2017kfm,Chen:2017sci,Agaev:2017lip,Wang:2017zjw}.
In 2018, the LHCb collaboration announced the observation of a
new bottom baryon $\Xi_b(6227)^-$ in both
$\Lambda_b K^-$ and $\Xi_b^0\pi^-$ decay modes~\cite{Aaij:2018yqz} and found two new resonances $\Sigma_b(6097)^{\pm}$
in the $\Lambda_b \pi^{\pm}$ channels~\cite{Aaij:2018tnn}. Both the $\Sigma_b(6097)^{\pm}$ and $\Xi_b(6227)^-$ states favor $P$ wave states with spin-parity $J^P = 3/2^-$ or $J^P = 5/2^-$\cite{Yang:2018lzg,Aliev:2018vye,Cui:2019dzj,Wang:2018fjm,Chen:2018orb,Chen:2018vuc}.
Recently, the LHCb collaboration announced the observation of two
new bottom baryons $\Lambda_b(6146)$ and $\Lambda_b(6152)$ in the
$\Lambda_b\pi^+\pi^-$ decay mode~\cite{Aaij:2019amv}.  The measured masses and widths of the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ are presented as follows,
\begin{eqnarray}
m[\Lambda_b(6146)] = 6146.17\pm0.33\pm0.22\pm0.16~\rm{MeV},
\end{eqnarray}
\begin{eqnarray}
\Gamma[\Lambda_b(6146)] = 2.9\pm1.3\pm0.3~\rm{MeV},
\end{eqnarray}
\begin{eqnarray}
m[\Lambda_b(6152)] = 6152.51\pm0.26\pm0.22\pm0.16~\rm{MeV},
\end{eqnarray}
\begin{eqnarray}
\Gamma[\Lambda_b(6152)] = 2.1\pm0.8\pm0.3~\rm{MeV}.
\end{eqnarray}
Considering their masses and small splitting, the LHCb Collaboration suggested that both of them may be clarified into the doublet of $\Lambda_b(1D)$ states in the quark model. It should be noticed that current experimental information does not exclude the two new resonances as the $I=0$ $\Sigma_b$ states, which has to be considered in the following discussions.

\begin{table*}[htp]
\begin{center}
\caption{\label{sp1}  Mass spectra of the $\Lambda_b$ up to $D$ wave and the $\Sigma_b$ up to $2S$ wave from
various quark models~\cite{Ebert:2011kk,Yoshida:2015tia,Roberts:2007ni,Capstick:1986bm,Chen:2014nyo} compared
with the data from the Particle Data Group~\cite{Tanabashi:2018oca}. The units are in MeV.}
\begin{tabular}{c|ccccccccccccccccccccccccccccccccccccccccccccc}\hline\hline
&\multicolumn{6}{c}{$\underline{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Lambda_b~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}$}\\
State            &~~~~\cite{Ebert:2011kk} &~~~~~~~~~~~~\cite{Yoshida:2015tia} &~~~~~~~~~~~~\cite{Roberts:2007ni} &~~~~~~~~~~~~\cite{Capstick:1986bm}   &~~~~~~~~~~~~\cite{Chen:2014nyo} &~~~~~~~~~~~~PDG\cite{Tanabashi:2018oca}\\ \hline
$|J^P=\frac{1}{2}^+,0 \rangle$        &~~~~5620      &~~~~~~~~~~~~5618   &~~~~~~~~~~~~5612  &~~~~~~~~~~~~5585    &~~~~~~~~~~~~5619 &~~~~~~~~~~~~5620 \\
$|J^P=\frac{1}{2}^-,1\rangle$         &~~~~5930      &~~~~~~~~~~~~5938   &~~~~~~~~~~~~5939  &~~~~~~~~~~~~5912    &~~~~~~~~~~~~5911 &~~~~~~~~~~~~5912\\
$|J^P=\frac{3}{2}^-,1\rangle$         &~~~~5942      &~~~~~~~~~~~~5939   &~~~~~~~~~~~~5941  &~~~~~~~~~~~~5920    &~~~~~~~~~~~~5920 &~~~~~~~~~~~~5920\\
$|J^P=\frac{1}{2}^+,0\rangle$         &~~~~6089      &~~~~~~~~~~~~6153   &~~~~~~~~~~~~6107  &~~~~~~~~~~~~6045       &~~~~~~~~~~~~$\cdot\cdot\cdot$     &~~~~~~~~~~~~? \\
$|J^P=\frac{3}{2}^+,2\rangle$         &~~~~6190      &~~~~~~~~~~~~6211   &~~~~~~~~~~~~6181  &~~~~~~~~~~~~6145    &~~~~~~~~~~~~6147 &~~~~~~~~~~~~ ?\\
$|J^P=\frac{5}{2}^+,2\rangle$         &~~~~6196      &~~~~~~~~~~~~6212   &~~~~~~~~~~~~6183  &~~~~~~~~~~~~6165    &~~~~~~~~~~~~6153 &~~~~~~~~~~~~?\\

\hline\hline
&\multicolumn{6}{c}{$\underline{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Sigma_b~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}$}\\
State        &~~~~\cite{Ebert:2011kk}       &~~~~~~~~~~~~ \cite{Yoshida:2015tia}  &~~~~~~~~~~~~\cite{Roberts:2007ni} &~~~~~~~~~~~~\cite{Capstick:1986bm}  &~~~~~~~~~~~~ &~~~~~~~~~~~~PDG~\cite{Tanabashi:2018oca}\\ \hline
$|J^P=\frac{1}{2}^+,1 \rangle$          &~~~~5808       &~~~~~~~~~~~~5823     &~~~~~~~~~~~~5833   &~~~~~~~~~~~~5795  &~~~~~~~~~~~~ &~~~~~~~~~~~~5811 \\
$|J^P=\frac{3}{2}^+, 1\rangle$          &~~~~5834       &~~~~~~~~~~~~5845     &~~~~~~~~~~~~5858   &~~~~~~~~~~~~5805  &~~~~~~~~~~~~      &~~~~~~~~~~~~5832 \\
$|J^P=\frac{1}{2}^-,0\rangle$           &~~~~6101       &~~~~~~~~~~~~6127     &~~~~~~~~~~~~6099   &~~~~~~~~~~~~6070  &~~~~~~~~~~~~      &~~~~~~~~~~~~?\\
$|J^P=\frac{1}{2}^-,1\rangle$           &~~~~6095       &~~~~~~~~~~~~6135     &~~~~~~~~~~~~6106   &~~~~~~~~~~~~6070  &~~~~~~~~~~~~        &~~~~~~~~~~~~ ?\\
$|J^P=\frac{3}{2}^-,1\rangle$           &~~~~6096       &~~~~~~~~~~~~6132     &~~~~~~~~~~~~6101   &~~~~~~~~~~~~6070  &~~~~~~~~~~~~            &~~~~~~~~~~~~?\\
$|J^P=\frac{3}{2}^-,2\rangle$           &~~~~6087       &~~~~~~~~~~~~6141     &~~~~~~~~~~~~6105   &~~~~~~~~~~~~6085  &~~~~~~~~~~~~           &~~~~~~~~~~~~ ?\\
$|J^P=\frac{5}{2}^-,2\rangle$           &~~~~6084       &~~~~~~~~~~~~6144     &~~~~~~~~~~~~6172   &~~~~~~~~~~~~6090  &~~~~~~~~~~~~            &~~~~~~~~~~~~?\\
$|J^P=\frac{1}{2}^+,1\rangle$           &~~~~6213       &~~~~~~~~~~~~6343     &~~~~~~~~~~~~6294   &~~~~~~~~~~~~6200  &~~~~~~~~~~~~          &~~~~~~~~~~~~?\\
$|J^P=\frac{3}{2}^+,1\rangle$           &~~~~6226       &~~~~~~~~~~~~         &~~~~~~~~~~~~6308   &~~~~~~~~~~~~6250  &~~~~~~~~~~~~           &~~~~~~~~~~~~ ?\\

\hline\hline
\end{tabular}
\end{center}
\end{table*}

Since the $\Lambda_b$ and $\Sigma_b$ baryons contains a heavy bottom quark and two light $u/d$ quarks, the low-lying internal excitations favor  excitations of the so-called $\lambda$-mode where the orbital excitation lies between the light
quarks and the heavy quark in a Jacobi coordinate. In the $\Lambda_b$ family,  the ground state $\Lambda_b$, and two low-lying orbitally excited states $\Lambda_b(5912)~1/2^-$ and $\Lambda_b(5920)~3/2^-$,  have been well established \cite{Capstick:1986bm,Ebert:2007nw,Roberts:2007ni,Ebert:2011kk,Chen:2014nyo,Garcilazo:2007eh,Karliner:2015ema,
Mao:2015gya,Thakkar:2016dna,Wang:2017kfr}.  According to the mass spectrum predicted in various models~\cite{Ebert:2007nw,Roberts:2007ni,Ebert:2011kk,Chen:2014nyo,Mao:2015gya,Capstick:2000qj,
Ebert:2005xj,Yoshida:2015tia,Valcarce:2008dr,Karliner:2008sv,Chen:2016phw}, the two newly observed resonances $\Lambda_b(6146)$ and $\Lambda_b(6152)$, lie in the mass region of the $\lambda$-mode $\Lambda_b(1D,2S)$ states and the neutral
excited $\Sigma_b(2S)$ states.

To identify the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ resonances in the quark model, their strong decay behaviors were studied
with the $^3P_0$ model in Ref.~\cite{Liang:2019aag}.
The results indicated that the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ can be assigned as the $\Lambda_b(1D)$ doublet with $J^P=5/2^+$ and $J^P=3/2^+$, respectively. In our previous works~\cite{Wang:2017kfr,Wang:2018fjm},  we have studied the strong decays of the singly bottom baryons $\Sigma_b(1P)$ within the chiral quark model.  We find that the $\Sigma_b(6097)$ favors the light
spin $j = 2$ states with spin-parity numbers $J^P=3/2^-$ or $J^P=5/2^-$. If $\Sigma_b(6097)$ corresponds to the $|J^P=\frac{3}{2}^-,2\rangle$ or  $|J^P=\frac{5}{2}^-,2\rangle$ state, the typical mass for the  $\Sigma_b(1P)$ states should be around 6090 MeV since the mass splitting between them is within 10 MeV according to quark model predictions~\cite{Ebert:2011kk}. Thus, one should exclude the $\Sigma_b(6146)$ and $\Sigma_b(6152)$ resonances as the $\Sigma_b(1P)$ states for the masses of $\Sigma_b(6146)$ and $\Sigma_b(6152)$ are obviously larger than the quark model predictions.

In previous works, the decays of the $P$- and $D$-wave singly heavy baryons have been studied within the chiral quark model~\cite{Yao:2018jmc,Wang:2017kfr}. Since the newly observed $\Lambda_b(6146)$ and $\Lambda_b(6152)$ resonances may favor the $D$-wave $\Lambda_b$ states, to confirm this assignment, in present work we revisit the strong decays of the $D$-wave $\Lambda_b$ states by adopting the measured masses. Moreover, the $2S$-wave $\Lambda_b$ and $\Sigma_b$ states were not investigated in the frame work of chiral quark model. Hence, in this work, as a supplement of Refs.~\cite{Yao:2018jmc,Wang:2017kfr}, we employ the chiral quark model to study the strong decays with emission of one light
pseudoscalar meson for the low-lying $\Lambda_b(1D,2S)$ and $\Sigma_b(2S)$ states.
Combining the masses, total widths and decay modes, our results suggest that the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ states may favor the $\Lambda_b(1D)$ $|J^P = \frac{5}{2}^+,2 \rangle$ and $|J^P = \frac{3}{2}^+,2 \rangle$ states, respectively.

This paper is organized as follows. The spectrum and notations are presented in Sec.~\ref{spectrum}. The chiral quark model are briefly introduced in Sec.~\ref{model}. The  strong decays of the low-lying $\Lambda_b(1D,2S)$ and $\Sigma_b(2S)$ states are estimated in Sec.~\ref{results}. A short summary is presented in the last section.

\section{SPECTROSCOPY }\label{spectrum}

\begin{table}[htp]
\begin{center}
\caption{\label{JJ1}  The classifications of the low-lying $1P$ and $2S$-wave states belonging to $\mathbf{6}_F$ in the $j-j$ coupling scheme.}
\begin{tabular}{p{2.0cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm} |}\hline\hline
$|J^P,j \rangle$      &$J^P$            &$j$  &$n_{\rho}$ &$\ell_{\rho}$  &$n_{\lambda}$  &$\ell_{\lambda}$    &$L$   &$s_{\rho}$    &$s_{Q}$                     \\ \hline
$|J^P=\frac{1}{2}^-,0\rangle$   &$\frac{1}{2}^-$  & 0    & 0    & 0      & 0          &1          &1       &1         &$\frac{1}{2}$         \\
$|J^P=\frac{1}{2}^-,1\rangle$   &$\frac{3}{2}^-$  & 1    & 0    & 0      & 0          &1          &1       &1         &$\frac{1}{2}$              \\
$|J^P=\frac{3}{2}^-,1\rangle$   &$\frac{1}{2}^-$  & 1    & 0    & 0      & 0          &1          &1       &1         &$\frac{1}{2}$           \\
$|J^P=\frac{3}{2}^-,2\rangle$   &$\frac{3}{2}^-$  & 2    & 0    & 0      & 0          &1          &1       &1         &$\frac{1}{2}$            \\
$|J^P=\frac{5}{2}^-,2\rangle$   &$\frac{5}{2}^-$  & 2    & 0    & 0      & 0          &1          &1       &1         &$\frac{1}{2}$            \\
$|J^P=\frac{1}{2}^+,1\rangle$   &$\frac{1}{2}^+$  & 1    & 0    & 0      & 1          &0          &0       &1         &$\frac{1}{2}$            \\
$|J^P=\frac{3}{2}^+,1\rangle$   &$\frac{3}{2}^+$  & 1    & 0    & 0      & 1          &0          &0       &1         &$\frac{1}{2}$            \\

\hline\hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[htp]
\begin{center}
\caption{\label{JJ2}  The classifications of the low-lying $1P$, $1D$ and $2S$-wave states belonging to $\mathbf{\bar{3}}_F$ in the $j-j$ coupling scheme.}
\begin{tabular}{p{2.0cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm}p{0.5cm} }\hline\hline
$|J^P,j \rangle$      &$J^P$            &$j$  &$n_{\rho}$ &$\ell_{\rho}$  &$n_{\lambda}$  &$\ell_{\lambda}$    &$L$   &$s_{\rho}$    &$s_{Q}$                     \\ \hline
$|J^P=\frac{1}{2}^-,1\rangle$   &$\frac{1}{2}^-$  & 0    &0    & 0    &0            &1          &1       &1         &$\frac{1}{2}$         \\
$|J^P=\frac{3}{2}^-,1\rangle$   &$\frac{3}{2}^-$  & 1    &0    & 0    &0            &1          &1       &0         &$\frac{1}{2}$ \\
$|J^P=\frac{1}{2}^+,0\rangle$   &$\frac{1}{2}^+$  & 0    &0    & 0    &1            &0          &0       &0         &$\frac{1}{2}$\\
$|J^P=\frac{3}{2}^+,2\rangle$   &$\frac{3}{2}^+$  & 2    &0    & 0    &0            &2          &2       &0         &$\frac{1}{2}$\\
$|J^P=\frac{5}{2}^+,2\rangle$   &$\frac{3}{2}^+$  & 2    &0    & 0    &0            &2          &2       &0         &$\frac{1}{2}$\\
\hline\hline
\end{tabular}
\end{center}
\end{table}

The heavy baryon containing a heavy quark violates
the SU(4) symmetry. However, the SU(3) symmetry between the other two
light quarks ($u$, $d$, or $s$) is approximately kept. According to
the symmetry, the heavy baryons containing a single heavy quark
belong to two different SU(3) flavor representations: the symmetric sextet $\mathbf{6}_F$ and
antisymmetric antitriplet $\bar{\mathbf{3}}_F$.
In the singly bottom baryons, $\Lambda_b$ and $\Xi_b^{0,-}$ belonging to $\bar{\mathbf{3}}_F$ representation and the $\Sigma_b^{-,0,+}$, $\Xi_b'^{0,-}$ and $\Omega_b$ form a $\mathbf{6}_F$ representation.

It should be pointed out that  the quark model states that we obtain for the
baryons containing a single heavy quark respect the dictates of the heavy quark effective theory (HQET).
In the heavy quark effective
theory description,
the total angular momentum of the two light quarks $\mathbf{j}=\mathbf{L}+\mathbf{s}_{\rho}$ is conserved and coupled to the spin of a heavy quark with spin $\mathbf{s}_Q=1/2$~\cite{Cheng:2015iom,Roberts:2007ni}. The total angular
momentum can take the values $\mathbf{J}=\mathbf{j}+\mathbf{s}_Q$.
In the heavy-quark
symmetry limit, the quark model states may more favor the
$j-j$ coupling scheme
\begin{eqnarray}
|J^P,j\rangle = |\{[(\ell_\rho \ell_\lambda)_L s_{\rho}]_js_Q\}_{J^P}\rangle.
\end{eqnarray}
The $\Sigma_b$ $1P$- and $2S$-wave states belonging to $\mathbf{6}_F$ and the $\Lambda_b$ $1P$-, $2S$- and $1D$-wave states belonging to $\bar{\mathbf{3}}_F$ in the $j-j$ coupling scheme and
their corresponding quantum numbers have been collected in Table~\ref{JJ1} and \ref{JJ2}.

\section{ Strong decay  }\label{model}

In this work,  strong decays of the  singly heavy baryons with emission of one light pseudoscalar meson, i.e. $\pi$, $K$ and $\eta$, are studied within chiral quark model~\cite{Manohar:1983md}.
In this model, the light pseudoscalar mesons
are treated as point-like Goldstone boson.
This model has been successfully applied to study the strong decays of heavy-light mesons
and strange and singly heavy baryons~\cite{Zhong:2008kd,Zhong:2010vq,Zhong:2009sk,
Zhong:2007gp,Liu:2012sj,Xiao:2013xi,Nagahiro:2016nsx,Wang:2017hej,Xiao:2014ura,Xiao:2017udy,Yao:2018jmc}.
The effective quark-pseudoscalar-meson interactions  at low energies
can be described by the simple chiral Lagrangian~\cite{Zhong:2008kd,Zhong:2010vq,Zhong:2009sk,
Zhong:2007gp,Liu:2012sj,Xiao:2013xi,Wang:2017hej,Xiao:2014ura}:
\begin{equation}\label{coup}
H_{m}=\sum_j
\frac{1}{f_m}\bar{\psi}_j\gamma^{j}_{\mu}\gamma^{j}_{5}\psi_j\partial^{\mu}\phi_m,
\end{equation}
where $\psi_j$ represents the $j$th quark field in the hadron;
$\phi_m$ is the pseudoscalar meson field, $f_m$ is the
pseudoscalar meson decay constant.

To match the nonrelativistic harmonic oscillator wave functions
in this work, one should adopt the quark-pseudoscalar-meson interactions in
the nonrelativistic form~\cite{Zhong:2008kd,Zhong:2010vq,Zhong:2009sk,
Zhong:2007gp,Liu:2012sj,Xiao:2013xi,Wang:2017hej,Xiao:2014ura}:
\begin{eqnarray}\label{ccpk}
H_{m}^{nr}=\sum_j\left[ \mathcal{G} \vsig_j \cdot \textbf{q}
+h \vsig_j\cdot \textbf{p}_j\right]I_j
e^{-i\mathbf{q}\cdot \mathbf{r}_j},
\end{eqnarray}
with $\mathcal{G}\equiv -(1+\frac{\omega_m}{E_f+M_f})$ and $h\equiv \frac{\omega_m}{2\mu_q}$.
In the above equation, $\omega_m$ and $\mathbf{q}$ are the energy and three momenta of the emitted light meson,
respectively;  $\mu_q$ stand for a reduced mass given by $1/\mu_q=1/m_j+1/m'_j$ with
$m_j$ and $m'_j$ for the masses of the $j$th quark in the initial and
final hadrons, respectively; $\vsig_j$ and $\textbf{p}_j$ are the
Pauli spin vector and internal momentum operator for the $j$th quark of the initial hadron;
and $I_j$ is the isospin operator associated with the pseudoscalar meson.

For a light pseudoscalar meson emission in a strong decay process,
the partial decay width can be calculated with~\cite{Zhong:2008kd, Zhong:2007gp}
\begin{equation}\label{dww}
\Gamma_m=\left(\frac{\delta}{f_m}\right)^2\frac{(E_f+M_f)|\mathbf{q}|}{4\pi
M_i(2J_i+1)} \sum_{J_{fz},J_{iz}}|\mathcal{M}_{J_{fz},J_{iz}}|^2 ,
\end{equation}
where $\mathcal{M}_{J_{fz},J_{iz}}$  corresponds to
the strong  amplitudes.
The quantum numbers $J_{iz}$ and $J_{fz}$ stand for the third components of the total
angular momenta of the initial and final heavy baryons,
respectively. $M_i$ is the mass of the initial heavy baryon. $E_f$ and $M_f$ are the energy and mass of the final heavy baryon.
$\delta$ as a global parameter accounts for the
strength of the quark-meson couplings.  It has been determined in our previous work of the strong
decays of the charmed baryons and heavy-light mesons
\cite{Zhong:2007gp,Zhong:2008kd}. Here, we fix its value the same as
that in Refs.~\cite{Zhong:2008kd,Zhong:2007gp}, i.e. $\delta=0.557$.

In the calculation, the standard quark model parameters are
adopted. Namely, we set $m_u=m_d=330$ MeV and $m_b=5000$ MeV for the constituent quark masses.
The harmonic oscillator parameter $\alpha_{\rho}$ in the wave function
$\psi^n_{lm}=R_{nl}Y_{lm}$ for $uu/ud/dd$ diquark systems is taken as $\alpha_{\rho}=400$ MeV. Another harmonic oscillator parameter $\alpha_{\lambda}$ can
be related to $\alpha_{\rho}$ with the relation $\alpha^2_\lambda=\sqrt{3m_b/(2m+m_b)}\alpha^2_\rho$. The decay
constant for $\pi$  meson is taken as $f_\pi = 132$ MeV. The masses of the well-established hadrons used in the calculations are taken from the Particle Data Group (PDG)~\cite{Tanabashi:2018oca}, and the masses of the undiscovered initial states adopt  from the predictions in Refs.~\cite{Ebert:2011kk,Capstick:1986bm}.

\section{Results and discussions}\label{results}

\begin{table}[htb]
\begin{center}
\caption{ \label{Lamb1D} Partial widths (MeV) of strong decays for the $\lambda$-mode $1D$-wave   $\Lambda_b$ baryons.}
%\footnotesize
\begin{tabular}{c|ccccc}
\hline\hline
 \multirow{2}{*}{Decay mode}        &\multicolumn{2}{c}{$\underline{~~~~~~|J^P = \frac{3}{2}^+,2 \rangle~~~~~~}$} &\multicolumn{2}{c}{$\underline{~~~~~~|J^P = \frac{5}{2}^+,2 \rangle~~~~~~}$} \\
             &~~ $\Lambda_b(6146)$         &$\Lambda_b(6152)$    &~~ $\Lambda_b(6146)$         &$\Lambda_b(6152)$ \\ \hline
  $\Sigma_b\pi$                                &4.41    &4.67     &0.64   &0.73  \\
 $\Sigma_b^*\pi$                               &1.26    &1.41     &4.26   &4.60  \\
Sum                                            &5.67    &6.08     &4.90   &5.33 \\\hline
\hline
\end{tabular}
\end{center}
\end{table}

%Our results are listed in Table.\ref{Lamb2S}

\subsection{$\Lambda_b(1D)$}

In the $\Lambda_b$ family,  there are two $\lambda$-mode $1D$-wave
excitations $|J^P = \frac{3}{2}^+,2 \rangle$
and $|J^P = \frac{5}{2}^+,2 \rangle$ according to the classification of quark models.
The masses for the $\lambda$-mode $1D$-wave $\Lambda_b$
excitations are predicted to be $\sim 6.1$-$6.2$ GeV (see Table~\ref{sp1}).
The measured masses of the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ indicate that they may be  good candidates of the $\lambda$-mode $1D$-wave excitations.

Our calculations are presented in Table~\ref{Lamb1D}. The $\Lambda_b(6146)$ resonance is most likely to be the $J^P = 5/2^+$ $|J^P = \frac{5}{2}^+,2 \rangle$ state. Assigning the $\Lambda_b(6146)$ as $|J^P = \frac{5}{2}^+,2 \rangle$,  one can find that it has a narrow width of $\sim$ 5 MeV, and dominantly decays into $\Sigma_b^*\pi$. The partial widths into $\Sigma_b\pi$ and $\Sigma_b^*\pi$ channels are predicted to be
\begin{eqnarray}\label{1DL}
\Gamma[\Lambda_b(6146) \rightarrow \Sigma_b\pi]\simeq 0.64 \ \mathrm{MeV}, \\
\Gamma[\Lambda_b(6146) \rightarrow \Sigma_b^*\pi]\simeq 4.26 \ \mathrm{MeV}.
\end{eqnarray}
Both the decay width and decay mode are consistent with the observations of $\Lambda_b(6146)$  considering the model uncertainties. This conclusion is consistent with that of the $^3P_0$ model~\cite{Liang:2019aag}.
To further confirm the nature of the $\Lambda_b(6146)$ resonance,  the partial width ratio between $\Sigma_b\pi$ and $\Sigma_b^*\pi$ channels,
\begin{eqnarray}\label{aa}
\frac{\Gamma(\Sigma_b\pi)}{\Gamma(\Sigma_b^*\pi)}\simeq 0.15,
\end{eqnarray}
is worth to observing in future experiments.

On the other hand, the  $\Lambda_b(6152)$ resonance most likely corresponds to the $\lambda$-mode $1D$-wave $\Lambda_b$ excitation $|J^P = \frac{3}{2}^+,2 \rangle$. When we assign the $\Lambda_b(6152)$ as $|J^P = \frac{3}{2}^+,2 \rangle$, it has a narrow width of $\sim$ 6 MeV, and dominantly decays into $\Sigma_b\pi$ and $\Sigma_b^*\pi$. We predicted partial widths
\begin{eqnarray}\label{1DL}
\Gamma[\Lambda_b(6152) \rightarrow \Sigma_b\pi]\simeq 4.67 \ \mathrm{MeV}, \\
\Gamma[\Lambda_b(6152) \rightarrow \Sigma_b^*\pi]\simeq 1.41 \ \mathrm{MeV},
\end{eqnarray}
which are in agreement with the the $^3P_0$ model predictions $\Gamma[\Lambda_b(6152) \rightarrow \Sigma_b\pi]\simeq 5.31 $ MeV and $\Gamma[\Lambda_b(6152) \rightarrow \Sigma_b\pi]\simeq 0.87 $ MeV~\cite{Liang:2019aag}. The partial width ratio between $\Sigma_b\pi$ and $\Sigma_b^*\pi$ channels is predicted to be
\begin{eqnarray}\label{aa}
\frac{\Gamma(\Sigma_b\pi)}{\Gamma(\Sigma_b^*\pi)}\simeq 3.3.
\end{eqnarray}
This ratio might be crucial to test the nature of $\Lambda_b(6152)$, which is suggested to be measured in future experiments.

Finally, it should be mentioned that the $\Lambda_c(2860)3/2^+$ and $\Lambda_c(2880)5/2^+$ are
often assigned to $1D$ doublet in the $\Lambda_c$ family~\cite{Chen:2016iyi,Chen:2017aqm,Yao:2018jmc}. It indicates that the mass of the $J^P=3/2^+$ $D$-wave state should be smaller than that of the $J^P=5/2^+$ state. Then, if assigning $\Lambda_b(6146)$ and $\Lambda_b(6152)$ to
be $J^P=5/2^+$ and $J^P=3/2^+$ $D$-wave states, respectively, one should face a serious problem of mass reverse.
Whether the structure in the $\Lambda_b\pi^+\pi^-$ spectrum correspond to two states or one state should be further
clarified with more observations in future experiments.

\subsection{$\Lambda_b(2S)$}

\begin{table}[htb]
\begin{center}
\caption{ \label{Lamb2S} Partial widths (MeV) of strong decays for the $\lambda$-mode $2S$-wave $\Lambda_b$ baryon.}
%\footnotesize
\begin{tabular}{c|ccccc}
\hline\hline
 \multirow{2}{*}{Decay mode}        &\multicolumn{2}{c}{$\underline{~~~~~~|J^P = \frac{1}{2}^+,0 \rangle~~~~~~}$} \\
             &~~ $\Sigma_b(6146)$         &$\Sigma_b(6152)$     \\ \hline
 $\Sigma_b\pi$                                       &0.34                    &0.32                    \\
 $\Sigma_b^*\pi$                                     &1.72                    &1.77   \\
Sum                                                  &2.06                    &2.09  \\
\hline
\hline
\end{tabular}
\end{center}
\end{table}

In the $\Lambda_b$ family, there is only one $\lambda$-mode $2S$-wave
excitation $|J^P = \frac{1}{2}^+,0 \rangle$ according to the quark model classification.
The mass for the $\lambda$-mode $2S$-wave $\Lambda_b$
excitation is predicted to be $\sim 6.1$ GeV in various quark models (see Table~\ref{sp1}).
According to the predicted masses in Ref.~\cite{Yoshida:2015tia},
the measured mass of the $\Lambda_b(6146)$ or $\Lambda_b(6152)$ indicates that it might be a
good candidate of the $\lambda$-mode $2S$-wave excitation.
Considering $\Lambda_b(6146)$ or $\Lambda_b(6152)$ as the $2S$-wave state,
the strong decay properties are studied, our results are listed in Table \ref{Lamb2S}.
It is found that if assigning $\Lambda_b(6146)$ to the $2S$-wave state,
both the total width $\Gamma\simeq 2$ MeV and the dominant decay mode $\Sigma_b^*\pi$
predicted in theory are consistent with the observations. However, the other
resonance $\Lambda_b(6152)$ cannot be understood in the quark model. It cannot be
assigned to any $D$-wave states in the $\Lambda_b$ family, because the mass splitting between the $2S$-wave
and $D$-wave $\Lambda_b$ states is $\sim50-100$ MeV. The $D$-wave $\Sigma_b$ states should
be excluded as well for their typical mass is $\sim 6.3$ GeV~\cite{Capstick:1986bm,Yao:2018jmc}.
As a whole if we assign $\Lambda_b(6146)$ as the
$\Lambda_b(2S)$ state, the other state $\Lambda_b(6152)$ cannot be reasonably explained according to the classification of quark models and mass splitting~\cite{Yoshida:2015tia}.

Since the $\Lambda_b(2S)$ $|J^P=\frac{1}{2}^+,0 \rangle$ state may not favor $\Lambda_b(6146)$, in this work, we take its mass $M=6045$ MeV as predicted in Ref.~\cite{Capstick:1986bm}, and estimate the strong decay of $\Lambda_b(2S)$  into the $\Sigma_b\pi$ and $\Sigma_b^*\pi$ channels. It is found that $\Lambda_b(2S)$ mainly decays into $\Sigma_b\pi$ and $\Sigma_b^*\pi$ modes. The predicted partial widths are
\begin{eqnarray}\label{1DL}
\Gamma[\Sigma_b\pi ]\simeq 0.21 \ \mathrm{MeV},~~~~~~ \Gamma[\Sigma_b^*\pi ]\simeq 0.39 \ \mathrm{MeV},
\end{eqnarray}
and the corresponding  total decay width reads
\begin{eqnarray}\label{1DL}
\Gamma_{\mathrm{total}}\simeq 0.6 \ \mathrm{MeV}.
\end{eqnarray}
The $\Sigma_b^*\pi$ might be an ideal channel for searching for the $\lambda$-mode $2S$-wave $\Lambda_b$ state $|J^P = \frac{1}{2}^+,0 \rangle$ in future experiments.

\begin{table}[htb]
\begin{center}
\caption{ \label{sigem2S} Partial widths (MeV) of strong decays for the $\lambda$-mode $2S$-wave   $\Sigma_b$ baryons.}
%\footnotesize
\begin{tabular}{c|ccccc}
\hline\hline
 \multirow{2}{*}{Decay mode}        &\multicolumn{1}{c}{$\underline{~~~~~~|J^P = \frac{1}{2}^+,1 \rangle~~~~~~}$} &\multicolumn{1}{c}{$\underline{~~~~~~|J^P = \frac{3}{2}^+,1 \rangle~~~~~~}$} \\
             &~~ $\Sigma_b(6213)$            &~~ $\Sigma_b(6226)$         \\ \hline
 $\Lambda_b\pi$                      &3.82                       &4.72      \\
 $\Sigma_b\pi$                       &$5.32\times10^{-2}$        &$3.63\times10^{-3}$     \\
 $\Sigma_b^*\pi$                     &$6.90\times10^{-2}$        &$1.07\times10^{-2}$      \\
Sum                                  &3.94                        &4.73      \\
\hline
\hline
\end{tabular}
\end{center}
\end{table}

\subsection{$\Sigma_b(2S)$}

We also study the strong decay properties of  two $\lambda$-mode $2S$-wave
excitations $|J^P = \frac{1}{2}^+,1 \rangle$ and $|J^P = \frac{3}{2}^+,1 \rangle$ in the $\Sigma_b$ family according to the quark model classification. The masses for the $\lambda$-mode $2S$-wave $\Sigma_b$
excitations are predicted to be $6.2\sim 6.3 $ GeV in various models (see Table~\ref{sp1}).

To study the strong decay properties of the $2S$ wave $\Sigma_b$ excitations, we adopt the predicted masses in Ref.~\cite{Ebert:2011kk}. Our results are listed in Table \ref{sigem2S}. One can see that $|J^P = \frac{1}{2}^+,1 \rangle$ and $|J^P = \frac{3}{2}^+,1 \rangle$ dominantly decay into the $\Lambda_b\pi$ channel. Their partial widths are predicted to be
\begin{eqnarray}\label{1DL}
\Gamma[\Sigma_b(6213)\rightarrow\Lambda_b\pi ]\simeq 3.82 \ \mathrm{MeV},\\
\Gamma[\Sigma_b(6226)\rightarrow\Lambda_b\pi ]\simeq 4.72 \ \mathrm{MeV}.
\end{eqnarray}
The $\Lambda_b\pi$ might be an ideal channel to look for  these radial excitations in future experiments.

Finally, it should be mentioned that we do not consider the $\Lambda_b(6146)$ and $\Lambda_b(6152)$ as the $2S$ $\Sigma_b$ states,
for the mainly decay  modes and masses of $\Sigma_b(2S)$ predicted in the quark model are inconsistent with the observations.

%and dominantly decay mode

\section{Summary}\label{suma}

In this work, we study the strong decays of the low-lying $\lambda$-mode  $\Lambda_b(1D,2S)$ and $\Sigma_b(2S)$ states. Our results indicate that the newly observed $\Lambda_b(6146)$ might be assigned as the $\Lambda_b(1D)$ $|J^P = \frac{5}{2}^+,2 \rangle$ state, which dominantly decays into $\Sigma_b^*\pi$ channel. The $\Lambda_b(6152)$ seems to favor the $\Lambda_b(1D)$ $|J^P = \frac{3}{2}^+,2 \rangle$ states, with this assignment, its decay behaviors are dominated by the $\Sigma_b\pi$ and $\Sigma_b^*\pi$ channels, which are consistent with the experimental observations. However, if we assign $\Lambda_b(6146)$ and $\Lambda_b(6152)$ to the $J^P=5/2^+$ and $J^P=3/2^+$ $D$-wave states, respectively, one should face a serious problem of mass reverse. Whether the structure in the $\Lambda_b\pi^+\pi^-$ spectrum corresponds to two states or one state should be further clarified with more observations in future experiments. Moreover, we find that the $\Lambda_b(2S)$ $|J^P = \frac{1}{2}^+,0 \rangle$ state mainly decays into $\Sigma_b^*\pi$, and the $\Sigma_b(2S)$ $|J^P = \frac{1}{2}^+,1 \rangle$ and $|J^P = \frac{3}{2}^+,1 \rangle$ dominantly decay into $\Lambda_b\pi$ with narrow widths $\Gamma \simeq 3.82$ MeV and $\Gamma \simeq 4.72$ MeV, respectively. These theoretical predictions may provide helpful information for future experimental searches.

\section*{  Acknowledgments }

This work is supported, in part, by the National Natural Science Foundation of China under Grants No.~11775078,
No.~U1832173, and No.~11705056.
%\appendix
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%555

\bibliographystyle{unsrt}

\end{document}